Circular No. 5716

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

α ORIONIS

S. L. S. Yang, A. M. Larson, and A. W. Irwin, University of Victoria; and G. A. H. Walker and A. R. Walker, University of British Columbia, report: “Precise measurements of the relative radial velocities of α Ori, made at the Dominion Astrophysical Observatory using the hydrogen-fluoride absorption cell, confirm the rapid increase reported by Hatzes and Cochran on IAUC 5713. Between 1992 Oct. 26 and 1993 Jan. 19, the radial velocity increased by almost 2.8 km/s. During the same period, there is also a small ‘filling in’ of the Ca II triplet line at 866 nm. The stellar surface temperature, as inferred from the changing equivalent widths of lines, has also decreased by 565 ± 170 K, which is consistent with the ∼ 0.45 mag drop in the brightness as reported by Guinan et al. on IAUC 5708.”

A. K. Dupree, Center for Astrophysics; E. Guinan, Villanova University; and M. Smith, Computer Sciences Corporation and IUE Observatory, report: “Our monitoring program of α Ori with spectroscopy, photometry, and radial velocity measures indicates that the current deep light minimum and radial velocity changes (see IAUC 5708, 5713, and above) resemble the star’s behavior during 1988/1989. Fading of α Ori in the visual and ultraviolet was accompanied by local changes in the radial velocity of 6-7 km/s over a 3 to 5-month time period. The current behavior of the supergiant seems typical of variations observed with periods of ∼ 400 days since the intensive monitoring began in 1984. Photospheric radial velocity variations of ∼ 6 km/s in 5 months are not uncommon (see Smith et al. 1989, A.J. 98, 2233; Dupree et al. 1991, ASP Conf. Ser. 11, 468).”

NO X-RAY PULSAR IN N63A K. Dennerl and M. Kürster withdraw their claim of a pulsar in N63A

(cf. IAUC 5714), noting that “the pulsed emission was caused by an instrumental effect, specifically that the observed modulation is related to the spacecraft wobble. By coincidence, the center of the N63A image was periodically covered by the wire intersections within the PSPC support grid (a mesh of wires spaced by 2′.9) during the wobble motion, resulting in a maximum of shading, while no such obscurations of the comparison source occurred. We conclude that our previous interpretation that a pulsar is reponsible for the variability can no longer be sustained. We apologize for the inconvenience.”

1993 March 1 Daniel W. E. Green